Theoceangirl (the_ocean_girl):

Anybody mind helping me with a few geometry questions for some homework? I'll medal :)

MsBrains (ms-brains):

I'd love to try. Post them. :)

OpenStudy (anonymous):

I would like to help.

Theoceangirl (the_ocean_girl):

http://prntscr.com/arsvd7 http://prntscr.com/arsvld http://prntscr.com/arsvt4 http://prntscr.com/arsw26 I'll take whatever help I can get :) Thank you both

Theoceangirl (the_ocean_girl):

@Ms-Brains @Smart_Pug325

OpenStudy (youngstudier):

@jabez177

OpenStudy (anonymous):

First one is 5^2 + x^2 = 13^2. Pythagorean's Theorem works because the triangle is a right-triangle as shown by the right angle at the bisection of the short end of the triangle and the diameter line. So using Pythagoreans it's x^2 = 144, which means x = 12.

Theoceangirl (the_ocean_girl):

@texaschic101

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Using the Pythagoras theorem, we get from the triangle: \(5^2 + x^2 = 13^2\) \(x^2 = 169 - 25\) \(x^2 = 144\) \(x = 12\)